OpenStudy (anonymous):

-8=-3-6 squared 3 x^2

OpenStudy (anonymous):

\[-8=-3-6\sqrt[3]{x ^{2}}\]

OpenStudy (anonymous):

First simplify the problem by moving the -3 to the other side.

OpenStudy (anonymous):

so that makes it \[-5=-6\sqrt[3]{x ^{2}}\]??

OpenStudy (anonymous):

Now divide -5 by -6.

OpenStudy (anonymous):

And yes!

OpenStudy (anonymous):

so it's \[1.2\sqrt[3]{x ^{2}}\]

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

Wait isn't -.83

OpenStudy (anonymous):

\[8.333333333=\sqrt[3]{x ^{2}}\]

OpenStudy (anonymous):

No negative because the negatives cancelled each other out.

OpenStudy (anonymous):

Alrighty what's after this?

OpenStudy (anonymous):

Now cube or but each side to the power of 3.

OpenStudy (anonymous):

oh and sorry i made a typo it is .8333333 not 8.33333

OpenStudy (anonymous):

\[0.57\sqrt[9]{x ^{2}}\]??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now square root .578703037

OpenStudy (anonymous):

.75?

OpenStudy (anonymous):

Yes, to be exact .7607257743 and you can plug that in and it 7.999999 which is the closest it can get.

OpenStudy (anonymous):

Do you think you can help me with \[\sqrt{169x^{5}}\]using rational exponent?

OpenStudy (anonymous):

uhm...Im a bit rusty but i can try, what exactly do you want done?

OpenStudy (anonymous):

Thank you for your help. I have 45 pieces of homework to do some I understand some I don't and their due tomorrow "/

OpenStudy (anonymous):

Oh, thats bad:(

OpenStudy (anonymous):

thanks for the medal!

OpenStudy (anonymous):

Yes I know and If I don't finish I won't graduate :( && Your welcome.

OpenStudy (anonymous):

Well I know 169 squared root is 13 but I need to what to do with x^5?

OpenStudy (anonymous):

thats harsh!

OpenStudy (anonymous):

Uhm...Let me think...

OpenStudy (anonymous):

I have choices to pick from but two I know are not it do you think this one would be it? \[13x ^{5/2}\]

OpenStudy (anonymous):

http://www.regentsprep.org/Regents/math/algtrig/ATO1/FractionalExp.htm check this website out it might help

OpenStudy (anonymous):

Wait im doing it...

OpenStudy (anonymous):

Thanks I'm going to go with the one I wrote on top :)

OpenStudy (anonymous):

Wait what are the other options?

OpenStudy (anonymous):

its the one I just typed in or\[13x ^{2/5}\] I going with the one on top because I know 5 rooted is 2.5 or something like that and the this one 2/5 is .4

OpenStudy (anonymous):

okay, i dont know lol

OpenStudy (anonymous):

I'm going to hope it's right :)

OpenStudy (anonymous):

haha if you plug in 1 they are both 13 so maybe!

OpenStudy (anonymous):

I'll know here in a bit.

OpenStudy (anonymous):

okay, so whats next?

OpenStudy (anonymous):

(5x^4y^-3)^-2?

OpenStudy (anonymous):

that times its self....

OpenStudy (anonymous):

\[(5x^{4}y ^{-3)^{-2}}\]

OpenStudy (anonymous):

Would I multiply 5 by exponent -2?

OpenStudy (anonymous):

(5x^4y^-2)(5x^4y^-3) im not sure what to do with the negative...

OpenStudy (anonymous):

(5x^4y^-3)(5x^4y^-3) rather

OpenStudy (anonymous):

so y6/25x^8?

OpenStudy (anonymous):

Let me check to make sure...

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

Okay i got it now!

OpenStudy (anonymous):

\[2/(5x^4y^-3)\]

OpenStudy (anonymous):

Wait...

OpenStudy (anonymous):

\[1/(5x^4y^-3)^{2}\]

OpenStudy (anonymous):

Alrighty

OpenStudy (anonymous):

then on the bottom it would be (5x^4y^-3)(5x^4y^-3)

OpenStudy (anonymous):

Your a good math teacher:)

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

Thanks:$

OpenStudy (anonymous):

Yeah do you like math?

OpenStudy (anonymous):

Yea! Except geometry